# Faux Polyglot: A Study on Information Disparity in Multilingual Large Language Models

### v1

@@ 4.2 Generation
We measured the preference in generation in 5 languages for 6 models (3 model families) across 2 settings with 27 queries and 16 queries each. In total we queried each model 5375 times555(27 + 16) queries *5 query-languages *5 alternate document language* 5 original document language.
@@ 4.2.1 RQ1g: Preference for Native language
On average, across the two premises, the Native Language documents were the source for the answer 42% (SE = 0.54%) of times compared to 29.58% (SE = 0.5)for Foreign Language Documents while only 8.55% (SE = 0.30%) contains information from both the sources, and 19.8% (SE= 0.43%) the answer was from none of the sources. This shows that the models systematically prefer to generate answers from the document in the Native Language as that of the query, even though diverse facts and perspectives were presented in a different language. Breakdown by model is in Appx. LABEL:sec:_appendix-generation.
@@ 4.2.2 RQ2g: Impact of query type on language preference
Figure 6: The source reference proportions for various query types. There are four possible source references: none, both, Native Document, and Foreign Document. For differential queries, which require the model to choose between answers from Native and Foreign Documents, both sources perform comparably, with a slight preference for Native Documents. However, the model also frequently scores high for none. In exclusive queries, targeting facts unique to one context, Native and Foreign Documents again perform similarly, though the none score remains high. For contradiction queries, where information is contradictory in both contexts, the model shows a high score for Native Documents and both. In broad queries, the model scores higher for Native Documents, while Foreign Documents lag behind, and it struggles to synthesize content from both sides. In neutral and confirmatory queries, we see similar effects of slightly preferring native while the desired output (both) remains negligible.
Unlike retrieval, we observed that depending on the query, the model’s preference for the document varied (Figure 6). For factual information, we observed a preference for Native Language documents over Foreign Language documents (Broad: $M_{native}$ = 57.05%, SE = 6.39% vs $M_{foreign}$ = 25.38%, SE = 3.75%) in broad queries. Similarly, for specific-contradiction queries, native documents were preferred (Specific-Contradiction: $M_{native}$ = 35.36%, SE = 5.52% vs $M_{foreign}$ = 12.13%, SE = 1.23%). However, for specific-exclusive (Specific-Exclusive: $M_{native}$ = 41.48%, SE = 4.52% vs $M_{foreign}$ = 39.88%, SE = 4.31%) and specific-differential queries (Specific-Differential: $M_{native}$ = 28.72%, SE = 3.94% vs $M_{foreign}$ = 27.61%, SE = 4.52%), we did not observe a significant difference between native and foreign documents.
@@ 4.2.3 RQ3g: Second Language Preference
We saw a preference for higher resource language as the Second Language across different queries in different languages, as shown in figure 5. Overall, across all models and query languages, we saw that on average, en (M = 39.56%, SE = 4.72%) > de (M = 35.82% SE = 4.35%) > zh (M = 33.93% SE = 4.35%) > ar (M = 35.82% SE = 4.08%) > hi (M = 32.79% SE = 3.95%). This also reflects the language resource, where English is the highest resource, and Hindi is the lowest.
@@ 5 Discussion
Through a comprehensive evaluation of the retrieval and generation phases of current multilingual RAG systems, we found that multilingual LLMs exhibit a clear preference for the language of queries and documents in high-resource languages. Given the premise of multilingual LLMs, breaking the linguistic barriers and information cocoon for global users, our results highlight a significant drawback in providing information parity and motivate further research. Below, we discuss our main findings and their implications.
@@ 5.1 Diverse, multilingual documents are hidden in retrieval
Retrieving diverse perspectives is a crucial phase in RAG systems. We found that the retrieval systems disproportionately prefer Native Documents in retrieval. 68% of the top-10 documents were Native Documents, which implies that today, most multilingual documents do not even reach the context of the LLMs in the generation phase.
@@ 5.2 LLMs prefer facts in Native Languages in knowledge conflict
In the generation experiment, we provided the model with diverse multilingual texts. Our results found that LLMs prefer facts from Native Documents and rarely present facts from both Native and Foreign perspectives. Our results indicate the LLM-powered search system could create and intensify the information cocoon by only presenting facts in the information seeker’s Native Language.
@@ 5.3 Only if you knew what you don’t know
Our results also found that to generate results from Foreign Documents, the user must issue specific-differential or specific-exclusive queries (queries that specifically target exclusive information in the foreign document) to LLMs, whereas general queries like broad or confirmatory queries (general queries like tell me about event A) prefer Native documents. This, however, is difficult in a real-world setting where users often don’t know the topic they want to search for, and the queries are more exploratory in nature.
@@ 5.4 High-resource languages dominate the information landscape
During our evaluation, we asked about the second language preference, i.e., the preferred language, when there was no source of Native Language documents. We found that during the generation phase, LLMs preferred high-resource languages as their second language. This is worrying because, often, when querying about international topics, the information is not present in the Native Language of the query. For instance, in Figure 1, if we query about the India-China conflict in Arabic, although there are facts about the conflict in Hindi and Chinese, which are the languages of the countries involved in the conflict, the model chooses to present the English viewpoint only. This behavior reinforces the views from dominant languages and further marginalizes the non-dominant views.
In our generation evaluation, “Both” is considered ideal, where the model draws information from both documents and “None" is considered a failure, where the model can not generate correct responses. Our results showed that models’ average performance increases when the query language is English or the documents in context are English. Such results again highlight the linguistic divide in LLMs. Details are in Appx. LABEL:sec:_appendix-all-none-both.
6 Related Work
6.1 Language Disparity in Multilingual LLMs
There has been a rise in the multilingual capabilities of large language models (LLMs) Qin et al. (2024b) with models now supporting over 100 languages. This rise has been facilitated by increasing multilingual datasets Qin et al. (2024a); Singh et al. (2024); Sun and Duh (2020) and benchmarks Siddhant et al. (2020); Ahuja et al. (2023); Zhang et al. (2023); Liang et al. (2020); Ruder et al. (2021) evaluating a diverse set of multilingual capabilities where recent multilingual LLMs such as GPT-4o have achieved STOA performanceWatts et al. (2024). However, this progress often hides the inherent language disparity that exists in Multilingual LLMs Yu et al. (2022); Verma et al. (2022); Joshi et al. (2021); where pretraining data of models such as Llama and GPT-4 is still predominantly English. Such language disparity affects the multilingual capabilities of LLMs Li et al. (2024b). The effect of language disparity in information parity has not been studied previously, making our study unique. Furthermore, such evaluations lack context realism Liao and Xiao (2023).
6.2 Knowledge Conflicts
In real-world settings, multilingual interactions will be driven by conversational search, where the LLMs will interact with often conflicting and contradictory information across different languages. Recently, there have been studies exploring how such knowledge conflicts affect the LLMs responses Xu et al. (2024); Jin et al. (2024b); Kortukov et al. (2024); Chen et al. (2022). In multilingual settings, Dong et al. (2024); Jin et al. (2024a); Boughorbel and Hawasly (2023); Li et al. (2023) explore how queries in different languages and cultural contexts affect the LLMs generation and preference for answers. However, in today’s conversational search systems, there are two phases, information retrieval and generation, and how they individually react to knowledge conflict has not been studied yet. Moreover, less is known about how diverse types of queries influence the model’s response to such knowledge conflicts, which has not been studied yet. For instance, Sharma et al. (2024) showed that users issued confirmatory queries while interacting with controversial information where knowledge conflicts are common and in response, LLMs mirrored the bias of the query, how such interactions shape up in multilingual settings is also not been explored before. In this paper, we aim to answer these unanswered questions.
@@ 7 Conclusion
The recent rise of multilingual LLMs and RAG has played a pivotal part in the exponential growth of conversational search across diverse users with different backgrounds, cultures, and languages. Such multilingual conversational searches have the ability to break linguistic barriers and provide users with information consisting of diverse viewpoints. However, we find that instead of offering diverse viewpoints, current multilingual LLMs struggle with information disparity across different languages by 1) failing to retrieve diverse, multilingual texts, 2) bias towards higher resource languages over lower resource languages and 3) preferring to generate from Native Facts over Foreign Document facts while not being able to generate facts from both the documents. This failure of current Multilingual LLM models to offer consistent information across different querying styles and languages leads to linguistic information cocoons. These results concern information parity and can lead to polarization and hinder constructive communication across different communities.
@@ 8 Limitations
Another limitation is that we did not probe into the effect of pre-training knowledge. Pre-training data of models often have bias and based on that bias the models have a preference to generate certain outputs, how those existing preferences affect the RAG pipeline when dealing with conflicting information is another future direction. Moreover, our study only focuses on language differences instead of cultural differences. People in different cultures may speak the same language and have diverse narratives. Cultural differences during information foraging might add to the observed effect of linguistic preference. Finally, during our Evaluation, we only focused on one architecture of the RAG system, i.e., retrieval based on cosine similarity. However, there are other architectures, such as summarization, rerank, etc., that we have not evaluated.
@@ Acknowledgments
This work is partially supported by Cohere For AI Research Grant.
@@ Appendix A Appendix
A.1 Cost of Experiments
We had 27 factual + 16 opinion-based queries. For all queries, we had 2 contexts: alternate and original. We had all permutations and combinations of these contexts across the 5 target languages, making it 5$\times$5 contexts for each premise (festival and war). We query each context with their respective query pool across all languages. This makes the total number of requests (27 factual queries + 16 opinion-based queries) $\times$ 5 query-languages $\times$ 5 original context languages $\times$ 5 alternate context languages. Each input context was around 1000 tokens and 526 output tokens. This results in a total of 5,375,000 input tokens and 2,872,250 output tokens per model. The cost of GPT-4o is \$5.00 / 1M tokens for inputs \$15.00 /1M tokens for outputs. This made the total cost of the experiments in generation phase for OpenAI GPT-4o \$680.
A.2 Prompt used in Dataset creation

### v3

@@ 4.2 Generation
We measured the preference in generation in 5 languages for 6 models (3 model families) across 2 settings with 27 queries and 16 queries each. In total we queried each model 5375 times222(27 + 16) queries *5 query-languages *5 alternate document language* 5 original document language.
@@ 4.2.1 RQ1g: Preference for Native language
On average, across the two premises, answers were derived from Native Language documents 42% of the time (SE = 0.54%), compared to 29.58% (SE = 0.5%) for Foreign Language documents. Only 8.55% of answers (SE = 0.30%) contained information from both sources, and for 19.8% (SE = 0.43%), the answer came from neither source. This indicates that models systematically prefer generating answers from documents in the same language as the query, even when diverse facts and perspectives are available in a different language. Breakdown by model can be found in the Appendix LABEL:sec:_appendix-generation.
@@ 4.2.2 RQ2g: Impact of query type on language preference
Figure 6: There are four possible source references: None, Both, Native, and Foreign. For differential queries, prompting a choice between Native and Foreign Documents, both sources perform similarly, though there’s a slight preference for Native Documents. The model often scores high on None. For exclusive queries, focusing on context-specific facts, Native and Foreign perform alike, but None score is still high. For contradiction queries, with conflicting information, the model scores high on Native Documents and Both. In broad queries, the model favors Native Documents over Foreign Documents and struggles to combine content from both. In neutral and confirmatory queries, there is a minor bias toward Native, while the preferred output (Both) remains minimal.
Unlike retrieval, the preference for the document varied based on the query(Figure 6). For factual information, we observed a preference for Native Language documents over Foreign Language documents (Broad: $M_{native}$ = 57.05%, SE = 6.39% vs $M_{foreign}$ = 25.38%, SE = 3.75%) in broad queries. Similarly, for specific-contradiction queries, native documents were preferred (Specific-Contradiction: $M_{native}$ = 35.36%, SE = 5.52% vs $M_{foreign}$ = 12.13%, SE = 1.23%). However, specific-exclusive (Specific-Exclusive: $M_{native}$ = 41.48%, SE = 4.52% vs $M_{foreign}$ = 39.88%, SE = 4.31%) and specific-differential queries (Specific-Differential: $M_{native}$ = 28.72%, SE = 3.94% vs $M_{foreign}$ = 27.61%, SE = 4.52%), did not show differences between native and foreign documents.
@@ 4.2.3 RQ3g: Second Language Preference
We saw a preference for higher resource language as the Second Language across different queries in different languages, as shown in figure 5. Overall, across all models and query languages, we saw that on average, en (M = 49.28%, SE = 1.14) > de (M = 44.48% SE = 1.45) > zh (M = 42.05% SE = 1.07) > hi (M = 40.77% SE = 0.59) > ar (M = 39.77% SE = 1.15). This also represents the resource of the language where English is the highest resource and Hindi and Arabic are the lowest.
@@ 5 Discussion
Through a comprehensive evaluation of current multilingual LLM-based information systems, we found a clear preference for documents in the native language of queries and foreign documents in high-resource languages. Given the premise of multilingual LLMs, breaking the linguistic barriers and filter bubbles for global users, our results highlight a significant risk in providing information parity and motivate further research. Below, we discuss our main findings and their implications.
@@ 5.1 Diverse, multilingual documents are hidden in retrieval
Retrieving diverse perspectives is crucial in RAG pipelines. However, we observed that cosine-similarity-based retrieval systems disproportionately favor Native Documents, with 68% of the top-10 documents being native. This suggests that most multilingual documents fail to reach the LLM during generation.
@@ 5.2 LLMs prefer facts in Native Languages in knowledge conflict
During generation, we provided the model with diverse multilingual docs yet the LLMs preferred facts from Native documents and rarely presented facts from Foreign perspectives. This indicates that current conversational search could accelerate filter bubbles by only presenting facts in the Native language of the user query.
@@ 5.3 Only if you knew what you don’t know
Our results also indicated that to obtain information from Foreign Documents, users must issue queries that specifically target exclusive information contained within these documents. In contrast, general queries, such as broad or confirmatory, prefer Native Documents. This becomes challenging in real-world scenarios, where user queries are often exploratory, and users don’t know what to ask.
@@ 5.4 High-resource languages dominate the information landscape
While assessing foreign language preferences, particularly when Native Language documents were unavailable, we found that LLMs favored high-resource languages. It is a concerning trend as, for international issues, relevant information often isn’t available in the query’s Native Language, reinforcing dominant linguistic perspectives and sidelining marginalized ones. Additionally, our findings indicated that models’ average performance in generating results with diverse perspectives increased when queries were in English or when English documents were present. These outcomes further underline the linguistic divide in LLMs. Details are in Appendix LABEL:sec:_appendix-all-none-both.
5.5 Societal Implications
In the long term, the systematic preference for high-resource languages poses significant threats to minority perspectives in the information system. Empirical studies have demonstrated the adverse effects of artificial intelligence bias on human perceptionSharma et al. (2024); Jakesch et al. (2023); Blanco et al. (2023), thus highlighting the importance of creating an information ecosystem that stifles diverse viewpoints both within individual communities and across different cultural groups. Furthermore, the persistent failure to transcend linguistic divides risks the creation of filter bubbles, ultimately jeopardizing the common knowledge necessary for democratic stabilityFarrell and Schneier (2018). Moreover, concentrated power over such AI technologies poses substantial risks, as it enables a few individuals to manipulate information flows, facilitating mass persuasion and eroding the foundations of democratic discourse. Such concentration not only diminishes the credibility and trust users place in these systems but also exacerbates the spread of misinformation.
5.6 Navigating Faux Polyglots
ChatGPT has approximately 4.7 billion users worldwideDuarte (2025), underscoring the need for strategies for effective information seeking, especially given that three of the top five user bases are in non-English speaking countries. Faux Polyglots are a new challenge for information literacy in conversational search, we need to educate users to know the limitation of these systems and create scaffolds to support diverse information seeking. From a users perspective, the only feasible solution is to issue queries probing multiple perspectives in multiple languages. This raises the cost of information seeking exponentially and is infeasible and hence industry and community should create scaffolds like 1) aggregate responses from diverse personas, 2) retrieve documents by diversity in language and perspective and 3) issue warnings to users engaging in confirmatory queries.
@@ 6 Conclusion
The recent surge in multilingual large language models (LLMs) and Retrieval Augmented Generation (RAG) has significantly expanded conversational search across varied linguistic and cultural demographics. This technology promises to transcend linguistic barriers, offering users access to diverse perspectives. In this study, involving 13 multilingual LLMs, 170 documents, five languages, and seven query types, we observe that rather than facilitating a breadth of perspectives, current multilingual LLMs struggle with disparities across languages in the following ways: they 1) show a systematic preference for documents in the Native language of the query, 2) exhibit bias favoring higher-resource languages over lower-resource languages when generating responses without relevant documents in the native language of the query, 3) fail to generate responses from "Both" documents and, 4) show this systematic biases across different query-types. This failure to offer consistent information across different querying styles and languages leads to linguistic information cocoons. These issues raise concerns about information parity, potentially leading to polarization and impeding constructive communication across communities.
@@ 7 Limitations
In this study, we use carefully curated artificial tasks to demonstrate an important phenomenon. We modeled our creation on the different information conflicts that we saw in the coverage of controversial topics 2. However, there are trade-offs to using such synthetic data instead of empirical data. Aligning with tasks such as the Winograd Schema ChallengeLevesque et al. (2012), our method allowed us to isolate confounding factors such as model’s parametric knowledge.
We did not probe the effect of pre-training extensively (Appx. LABEL:appx:_generation-chinese). Pre-training data of models often have bias and based on that bias the models have a preference to generate certain outputs, how those existing preferences affect the RAG pipeline when dealing with conflicting information is another future direction. Moreover, our study only focuses on language differences instead of cultural differences. People in different cultures may speak the same language and have diverse narratives. Cultural differences during information foraging might add to the observed effect of linguistic preference. Finally, during our Evaluation, we only focused on one architecture of the RAG system, i.e., retrieval based on cosine similarity. However, there are other architectures, such as summarization, rerank, etc., that we have not evaluated.
8 Ethical Consideration
This paper explores the biases of the multilingual RAG system on topics that have conflicting perspectives. As our study shows, currently, RAG-based LLMs might cause information disparity and biases in their responses to various types of user queries. Our results highlight the harmful consequences of such models, such as information cocoons, and encourage users of such systems to exercise caution and we encourage the research community to provide scaffolds to users as suggested in Section 5.6. Unfortunately, currently none of the information ecosystems provide low-cost access to diverse information breaking language barriers but with multilingual LLMs we have the opportunity to develop systems that create common ground and lower polarization across these divisive topics.
@@ Acknowledgments
This work is partially supported by Cohere for AI Grant. Thanks to Sunil Sharma for the discussion on information-warfare motivating this project.
@@ Appendix A Appendix
A.1 Types of information conflicts
Our dataset supports two common information conflicts: facts and opinions. We considered the following scenarios for each type of query:
Discrepancies in facts:
  • Shared Facts: same facts shared across different languages.
Diverse opinions and perspectives:
A.2 Synthetic dataset setup
Core facts:
Extended story facts:
Final Documents:
Manipulation Check:
Multilingual Dataset
A.3 Cost of experiments
We had 27 factual + 16 opinion-based queries. For all queries, we had 2 contexts: alternate and original. We had all permutations and combinations of these contexts across the 5 target languages, making it 5$\times$5 contexts for each premise (festival and war). We query each context with their respective query pool across all languages. This makes the total number of requests (27 factual queries + 16 opinion-based queries) $\times$ 5 query-languages $\times$ 5 original context languages $\times$ 5 alternate context languages. Each input context was around 1000 tokens and 526 output tokens. This results in a total of 5,375,000 input tokens and 2,872,250 output tokens per model. The cost of GPT-4o is \$5.00 / 1M tokens for inputs \$15.00 /1M tokens for outputs. This made the total cost of the experiments in the generation phase for OpenAI GPT-4o \$680.
A.4 Prompt used in dataset creation